# Reliability entails input-selective contraction and regulation in excitable networks

### v1

@@ II-B Input-Selective Contraction of FitzHugh-Nagumo Models
For every $\mu>0$, system (1) is contractive on both $C_{\downarrow}(\mu)$ and $C^{\uparrow}(\mu)$ with respect to the metric (3) and with contraction rate $\lambda(\mu)=\sqrt{\min\{\mu,b\}}$.
@@
Moreover, $d$ satisfies $d(x_{a},x_{b})^{2}=(x_{a}-x_{b})^{\top}P(x_{a}-x_{b})$ with $P\coloneqq\frac{1}{2}\operatorname{diag}(1,\ \varepsilon^{-1})$. Then, for every $\mu>0$ and every $x\in C^{\uparrow}(\mu)\cup C_{\downarrow}(\mu)$, $J(x)^{\top}P+P^{\top}J(x)=\operatorname{diag}(1-v^{2},-b)\leq-\min\{\mu,b\}I$. The claim then follows from this latter inequality and from the convexity of $C^{\uparrow}(\mu)$ and $C_{\downarrow}(\mu)$ by means of well-known arguments based on Hadamard’s Lemma; see, e.g., [9, Lem 3.1, Thm. 3.9].
@@
Let $\alpha\coloneq e^{-\lambda(\mu)\Delta_{\rm c}(\mu)+\sqrt{2}(t_{1}-t_{0}-\Delta_{\rm c}(\mu))}$. If $\Delta_{\rm c}(\mu)>\sqrt{2}(\sqrt{2}+\lambda(\mu))^{-1}(t_{1}-t_{0})$, then $\alpha\in[0,1)$ and any two solutions in $\mathcal{S}(X_{0},u,t_{0})$ $\alpha$-contract from $t_{0}$ to $t_{1}$ with respect to the metric (3).
@@
Let $T_{\rm c}(\mu)$ denote the set of time instant in which all solutions in $\mathcal{S}(X_{0},u,t_{0})$ are altogether either in $C^{\uparrow}(\mu)$ or in $C_{\downarrow}(\mu)$. Pick any two solutions $x_{a},x_{b}\in\mathcal{S}(X_{0},u,t_{0})$. In view of Proposition 1, for every $\tau,s\in\mathbb{R}$ such that $(\tau,s)\subset T_{\rm c}(\mu)$, $d(x_{a}(s),x_{b}(s))\leq e^{-\lambda(\mu)(s-\tau)}d(x_{a}(\tau),x_{b}(\tau))$. Moreover, in view of (5), and since $v_{a}^{2}+v_{b}^{2}+v_{a}v_{b}\geq 0$ for all $v_{a},v_{b}\in\mathbb{R}$, for every $t\in[t_{0},t_{1}]\setminus T_{\rm c}(\mu)$ we obtain $\frac{\rm d}{{\rm d}t}d(x_{a}(t),x_{b}(t))^{2}\leq 2d(x_{a}(t),x_{b}(t))^{2}$. Grönwall’s Lemma, thus implies $d(x_{a}(t_{1}),x_{b}(t_{1}))\leq e^{-\lambda(\mu)\Delta_{\rm c}(\mu)+\sqrt{2}(t_{1}-t_{0}-\Delta_{\rm c}(\mu))}d(x_{a}(t_{0}),x_{b}(t_{0}))$, from which the results follows.
@@ IV Discussion, Implications, Conclusions
While in this LTI case we cannot talk about contraction in the strict sense, the observed behavior has all the elements of the input-selective contraction property of excitable systems studied in this paper. And, indeed, linear oscillators are the very heart of Francis and Wonham’s original internal model principle [17], and their proper embedding in the control loop is both necessary and sufficient for regulation in the LTI case. In addition, as proved in [20, Prop. 3], in the presence of contraction (incremental input-to-state stability), internal models made of LTI oscillators also guarantee robust rejection of harmonics at the embedded frequencies in general nonlinear systems. As sinusoids can be seen as a limit-case of events maximally concentrated in frequency and maximally spread in time – i.e., the presence of an harmonic – this ultimately creates a link to event regulation, which, in these terms, appears as a robust property also in classical nonlinear regulation.

### v2

@@ II-B Input-Selective Contraction of FitzHugh-Nagumo Models
For every $\mu>0$, system (1) is contractive on both $C_{\downarrow}(\mu)$ and $C^{\uparrow}(\mu)$ with respect to the metric (3) and with contraction rate $\lambda(\mu)=\min\{\mu,b\varepsilon\}$.
@@
Moreover, $d$ satisfies $d(x_{a},x_{b})^{2}=(x_{a}-x_{b})^{\top}P(x_{a}-x_{b})$ with $P\coloneqq\frac{1}{2}\operatorname{diag}(1,\ \varepsilon^{-1})$. Then, for every $\mu>0$ and every $x\in C^{\uparrow}(\mu)\cup C_{\downarrow}(\mu)$, $J(x)^{\top}P+P^{\top}J(x)=\operatorname{diag}(1-v^{2},-b)\leq-\operatorname{diag}(\mu,b)$. The claim then follows from this latter inequality and from the convexity of $C^{\uparrow}(\mu)$ and $C_{\downarrow}(\mu)$ by means of well-known arguments based on Hadamard’s Lemma; see, e.g., [6, Lem 3.1, Thm. 3.9].
@@
With $\lambda(\mu)=\min\{\mu,b\varepsilon\}$, let $\alpha\coloneq e^{-\lambda(\mu)\Delta_{\rm c}(\mu)-\Delta_{\rm c}(\mu)+t_{1}-t_{0}}$. If $\Delta_{\rm c}(\mu)>(t_{1}-t_{0})/(1+\lambda(\mu))$, then $\alpha\in[0,1)$ and any two solutions in $\mathcal{S}(X_{0},u,t_{0})$ $\alpha$-contract from $t_{0}$ to $t_{1}$ with respect to the metric (3).
@@
Let $T_{\rm c}(\mu)$ denote the set of time instant in which all solutions in $\mathcal{S}(X_{0},u,t_{0})$ are either all in $C^{\uparrow}(\mu)$ or all in $C_{\downarrow}(\mu)$. Pick any two solutions $x_{a},x_{b}\in\mathcal{S}(X_{0},u,t_{0})$. In view of Proposition 1, for every $\tau,s\in\mathbb{R}$ such that $(\tau,s)\subset T_{\rm c}(\mu)$, $d(x_{a}(s),x_{b}(s))\leq e^{-\lambda(\mu)(s-\tau)}d(x_{a}(\tau),x_{b}(\tau))$. Moreover, in view of (5), and since $v_{a}^{2}+v_{b}^{2}+v_{a}v_{b}\geq 0$ for all $v_{a},v_{b}\in\mathbb{R}$, for every $t\in[t_{0},t_{1}]\setminus T_{\rm c}(\mu)$ we obtain $\frac{\rm d}{{\rm d}t}d(x_{a}(t),x_{b}(t))^{2}\leq 2d(x_{a}(t),x_{b}(t))^{2}$. Grönwall’s Lemma, thus implies $d(x_{a}(t_{1}),x_{b}(t_{1}))\leq e^{-\lambda(\mu)\Delta_{\rm c}(\mu)+t_{1}-t_{0}-\Delta_{\rm c}(\mu)}d(x_{a}(t_{0}),x_{b}(t_{0}))$, from which the results follows.
@@ IV Discussion, Implications, Conclusions
While in this LTI case we cannot talk about contraction in the strict sense, the observed behavior has all the elements of the input-selective contraction property of excitable systems studied in this paper. And, indeed, linear oscillators are the very heart of Francis and Wonham’s original internal model principle [10], and their proper embedding in the control loop is both necessary and sufficient for regulation in the LTI case. In addition, as proved in [2, Prop. 3], in the presence of contraction (incremental input-to-state stability), internal models made of LTI oscillators also guarantee robust rejection of harmonics at the embedded frequencies in general nonlinear systems. As sinusoids can be seen as a limit-case of events maximally concentrated in frequency and maximally spread in time – i.e., the presence of an harmonic – this ultimately creates a link to event regulation, which, in these terms, appears as a robust property also in classical nonlinear regulation.